The area of a rectangle is 96ft square.
The width is 4 less than the length.
What is the perimeter?
I am having trouble coming up with a formula for questions similar to this.
L*W = 96, and L-4=W is all I got.

Answer by Alan3354(69443)

The area of a rectangle is 96ft square.
The width is 4 less than the length.
What is the perimeter?
I am having trouble coming up with a formula for questions similar to this.
L*W = 96, and L-4=W is all I got.
----------------
That's all you need.
Sub for W in the 1st eqn
L*W = 96
L*(L-4) = 96
L^2 - 4L - 96 = 0
(L - 12)*(L + 8) = 0
L = 12
--------
W = 8
----
You could have tried factors of 96 and gotten 8 by 12, too.
Perimeter = 40